In the Brown Hydrogenation reaction,

1. How is the zero-valent Pt catalyst produced? write the chemical equation

2. how is the hydrogen gas produced? Write the chemical equationNaBH2-HCI PECLA COH COZH CO2H CO2H cis-norbornene-5,6-endo- dicarboxylic acid mp=180-190 °C MW=182.17 cis-norbornane-5,6-endo
Answer #1

zero valent Pt produce during reduction of Pt(IV) by sodium borohydride consequently hydrogen gas is produced from BH3 in presence of HCl.

PtCl4 + 4NaBH4 = Pt(0) +4NaCl + 4BH3 +2H2

BH3 + 3HCl = BCl3 +3H2